OpenStudy (richyw):

I'm trying to plot a graph using mathematica. I would like to make a plot of the first few energy levels for the 1d particle in a box. so what I am basically trying to do is define a function of x, that depends on a parameter n, and I would like to plot that function with a few different parameters.

OpenStudy (richyw):

\[\psi (x)=\sqrt{\frac{2}{L}}\sin\left(\frac{n\pi x}{L}\right)\]where n=1,2,3...

OpenStudy (richyw):

So I want a plot from 0 to L, (the function is 0 elsewhere) that shows \(\psi(x)\) for like n=1, n=2 and n=3